OpenStudy (nathanjhw):

Given that f(x) is continuous, integral (-2,2) f(x) dx = 7, integral (0,4) f(x) dx =-3, and integral (-2,4) f(x) dx = 2. Then integral (0,2) f(x) dx = ?

OpenStudy (nathanjhw):

2 4 -3 -6 0

OpenStudy (zarkon):

\[\int\limits_{-2}^2f(x)dx+\int\limits_{0}^4f(x)dx-\int\limits_{-2}^4f(x)dx\]

OpenStudy (nathanjhw):

7 + (-3) - 2 = 2